OpenStudy (anonymous):

Write a vector equation of the line that passes through P(-1, 6) and is parallel to a (vector) = (3, 1)

OpenStudy (anonymous):

a. (x+7, y=8) = t(5, 1) b. (x-5, y-14) = t(8, 5) c. (x+1, y-6) = t(3, -1) d. (x+14, y+13) = t(-1, 8)

OpenStudy (anonymous):

@surjithayer @Blank  @TQKMB @iPwnBunnies

OpenStudy (blank ):

C. (x+1, y-6) = t(3, -1)

OpenStudy (anonymous):

r=a+mb a is the point through which the line passes and b is the direction of the line ,m is a scalar. \[r=-1 i+6 j+m(3i+1j)\]